OpenStudy (anonymous):

What is the simplified form of the ninth root of x times the ninth root of x times the ninth root of x times the ninth root of x? Need help!

OpenStudy (anonymous):

(x^(1/9))*(x^(1/9))*(x^(1/9))*(x^(1/9))

OpenStudy (anonymous):

\[\sqrt[9]{x}\cdot \sqrt[9]{x}\cdot \sqrt[9]{x}\cdot \sqrt[9]{x}= \sqrt[9]{x^4}=x^{\frac{4}{9}}\]

OpenStudy (anonymous):

what i said is same.....@pgpilot326

OpenStudy (anonymous):

really @kc_kennylau ?

OpenStudy (kc_kennylau):

lol

OpenStudy (anonymous):

for your question (x^(1/9))*(x^(1/9))*(x^(1/9))*(x^(1/9))...................................................(x^(1/9))*(x^(1/9))*(x^(1/9))*(x^(1/9)) @kc_kennylau

OpenStudy (kc_kennylau):

@skullpatrol calm down bro, just kidding

OpenStudy (kc_kennylau):

nope, relax bro, relax.

OpenStudy (kc_kennylau):

ok sorry won't do it again @skullpatrol